# On some practical challenges of conformal prediction

### v1

@@ 2 Answers to Questions (I)—(IX)
Answer to Question (V): No. (3) is not a necessary condition for $C_{\alpha}(X_{n+1},Z^{n})$ to be an interval.
Example 5.
$C_{\alpha}(X_{n+1},Z^{n})=(-\infty,a_{(k)}),$
Answer to Question (VI): Yes. Monotonicity of the plausibility function $\mathsf{pl}_{x_{n+1},z^{n}}$ is a necessary condition for $C_{\alpha}(X_{n+1},Z^{n})$ to be a one-sided interval.
Proof.
Answer to Question (VII): No. Monotonicity of $M$ is not a necessary condition for the exact determination of $C_{\alpha}(X_{n+1},Z^{n})$.
@@
Example 7.
Consider Example 5.
Example 8.
Consider Example 4.
Remark. We did not ask the converse of Question (IX), i.e., whether the monotonicity of $M$ implies that $C_{\alpha}(X_{n+1},Z^{n})$ can be determined exactly, because that question seems to be too broad to be well-defined. The next example illustrates this point.
@@ 3 Proposed strategy
If $\lfloor(n+1)\alpha\rfloor\leq 1$, then the $(1-\alpha)\%$ prediction region $C_{\alpha}(x;Z^{n})$ given by (1) is $\mathbb{R}$.
@@
We discuss two possible cases: (i) $\lfloor(n+1)\alpha\rfloor<1$ and (ii) $\lfloor(n+1)\alpha\rfloor=1$. We have $Y_{n+1}\in C_{\alpha}(X_{n+},Z^{n})$ if and only if $\mathsf{pl}_{X_{n+1},Z^{n}}(Y_{n+1})>\lfloor(n+1)\alpha\rfloor/(n+1)$ if and only if $\sum_{i=1}^{n+1}1_{\{\mu_{i}\geq\mu_{n+1}\}}>\lfloor(n+1)\alpha\rfloor$. Since $1_{\{\mu_{n+1}\geq\mu_{n+1}\}}=1$, we have $C_{\alpha}(X_{n+1},Z^{n})=\mathbb{R}$ in Case (i). In Case (ii), we have $1/(n+1)\leq\alpha<2/(n+1)$ and $t_{n}(\alpha)=1/(n+1)$. Since $\mu_{1},\ldots,\mu_{n+1}$ are exchangeable, $\mathsf{pl}_{X_{n+1},Z^{n}}(Y_{n+1})$ follows the discrete uniform distribution on the set $\{1/(n+1),2/(n+1),\ldots,n/(n+1),1\}$. Therefore,
$\mathsf{P}^{n+1}\{\mathsf{pl}_{X_{n+1},Z^{n}}(Y_{n+1})>t_{n}(\alpha)\}=\frac{n}{n+1}\geq 1-\alpha,\quad\text{for all $Y_{n+1}$}.$
It follows that $C_{\alpha}(X_{n+1},Z^{n})=\mathbb{R}$. ∎
Thus, if we want a nontrivial conformal prediction region (i.e., $C_{\alpha}(x;Z^{n})\neq\mathbb{R}$), we must require $\lfloor(n+1)\alpha\rfloor\geq 2$. For the remainder of this article, we let $r_{1}=\min\{n,\lfloor(n+1)(1-\alpha)\rfloor+1\}$, $r_{2}=(n+1)-\lfloor(n+1)(1-\alpha)\rfloor$, and $r_{3}=(n+1)-\lfloor(n+1)\alpha\rfloor$.
The negative answers to most questions in the previous section show that it is challenging to give a hard-and-fast rule for choosing a non-conformity measure so that the resulting conformal prediction region can be determined exactly and can be of the desired shape. In the extant literature, many existing statistical models, such as ordinary linear regression, ridge regression, lasso, and kernel density estimation, have been used to create non-conformity measures; see Vovk et al. (2005) and Lei et al. (2014), and references therein. However, barely any work discusses how to determine exactly the corresponding conformal prediction regions based on these non-conformity measures; the numerical examples based on these proposed non-conformity measures are mostly based on approximate determination of the corresponding conformal prediction regions. As pointed out in Section 2, the resulting regions have no provable finite-sample validity. Since conformal prediction is a model-free method, one does not have to use a complicated statistical model to create a non-conformity measure. There is no evidence, either theoretical or practical, that doing so would have any advantage. It is evident that if (3) holds, then we can address the three practical challenges simultaneously. There are numerous choices of the non-conformity measures that can lead to (3). Driven by the principle of parsimony, we prefer something simple, such as a linear polynomial in data. However, to obtain a two-sided/bounded prediction interval, we will need something other than (3). In fact, a multivariate polynomial of data with degree two suffices, as we will see below. Therefore, we propose the following non-conformity measure.
@@
$M(B,z)=(\beta_{2}y^{2}+\beta_{1}y)+\gamma\left[\sum_{j=1}^{p}x_{j}+\eta\sum_{i=1}^{n}\left(y_{i}-\sum_{j=1}^{p}x_{ij}\right)\right],$ (6)
where $\beta_{1},\beta_{2},\gamma$, and $\eta$ are constants to be chosen at the discretion of the data scientist.
The next three theorems show that this non-conformity measure has several advantages. First, the aforementioned three practical challenges will be addressed simultaneously. In particular, we can determine the resulting conformal prediction regions exactly. Second, it can generate conformal prediction intervals of three different shapes. Indeed, $\beta_{2},\beta_{1}$, and $\gamma$ control the shape of the conformal prediction regions. For example, if $\beta_{2}=0$ and $\beta_{1}=1$, the resulting conformal prediction region will be a one-sided interval; when $\beta_{2}\neq 0$ and $\beta_{1}=0$, the resulting conformal prediction region will be a bounded (two-sided) interval.
We do not use a higher-order polynomial because such a choice will involve more complicated computation, and will likely lead to a challenging case in determining the conformal prediction regions, similar to what we saw in Example 9. In the next three theorems, we take the coefficients of the second-order polynomial in (6) to be some simple numbers, such as $0,1$, and 1-1. Tedious algebra shows that using other coefficients will not provide any advantages.
@@
Suppose $0<\alpha<1$, $\lfloor(n+1)\alpha\rfloor\geq 2$, and the non-conformity measure is given by (6). If $\beta_{2}=0$, $\beta_{1}=1$, $\gamma=-1$, and $1+\eta>0$, then the $100(1-\alpha)\%$ conformal prediction region $C_{\alpha}(X_{n+1};Z^{n})$ is the one-sided interval $(-\infty,a_{(r_{1})})$, where $a_{i}=\sum_{j=1}^{p}(X_{(n+1)j}-X_{ij})+Y_{i}$ for $1\leq i\leq n$ and $a_{(k)}$ is the $k$-th ordered value of $a_{1},\ldots,a_{n}$.
@@
$\mu_{i}$ $=$ $M(Z^{n+1}\backslash\{Z_{i}\},Z_{i})=Y_{i}-\left[S_{i}+\eta\sum_{j=1,j\neq i}^{n+1}(Y_{j}-S_{j})\right],\ i=1,\ldots,n,$
$\mu_{n+1}$ $=$ $M(Z^{n},Z_{i})=Y_{n+1}-\left[S_{n+1}+\eta\sum_{j=1}^{n}(Y_{j}-S_{j})\right].$
@@
$S_{i}-Y_{i}+\eta\sum_{j=1}^{n}(Y_{j}-S_{j})-\eta(Y_{i}-S_{i})+\eta(Y_{n+1}-S_{n+1})\leq S_{n+1}-Y_{n+1}+\eta\sum_{j=1}^{n}(Y_{j}-S_{j}).$
Since $1+\eta>0$, the last display implies $\mu_{i}\geq\mu_{n+1}$ if and only if $Y_{n+1}\leq(S_{n+1}-S_{i})+Y_{i}$ for $1\leq i\leq n$. Therefore, the theorem follows from (1) and the definition of the plausibility function $\mathsf{pl}_{x_{n+1},z^{n}}$. ∎
@@
Suppose $0<\alpha<1$, $\lfloor(n+1)\alpha\rfloor\geq 2$, and the non-conformity measure is given by (6). If $\beta_{2}=0$, $\beta_{1}=-1$, $\gamma=1$, and $1+\eta<0$ and $\eta\neq-1$,, then the $100(1-\alpha)\%$ conformal prediction region $C_{\alpha}(X_{n+1};Z^{n})$ is the one-sided interval $(a_{(r_{2})},\infty)$, where $a_{i}=\sum_{j=1}^{p}(X_{(n+1)j}-X_{ij})+Y_{i}$ for $1\leq i\leq n$.
@@
Suppose $0<\alpha<1$, $\lfloor(n+1)\alpha\rfloor\geq 2$, and the non-conformity measure is given by (6). If $\beta_{2}=1$, $\beta_{1}=0$, $\gamma=-1$, and $\eta\geq\max_{1\leq i\leq n}\{2(\sqrt{\max\{0,c_{i}^{2}-d_{i}\}}-c_{i})\}$ where $c_{i}=Y_{i}+\sum_{j=1}^{p}(X_{(n+1)j}-X_{ij})$ and $d_{i}=Y_{i}^{2}+\sum_{j=1}^{p}(X_{(n+1)j}-X_{ij})$ for $1\leq i\leq n$, then the $100(1-\alpha)\%$ conformal prediction region $C_{\alpha}(X_{n+1};Z^{n})$ is the bounded interval $(a_{(r_{1})},b_{(r_{3})})$, where $a_{i}=-\sqrt{\eta^{2}/4+\eta c_{i}+d_{i}}-\eta/2$ and $b_{i}=\sqrt{\eta^{2}/4+\eta c_{i}+d_{i}}-\eta/2$ for $1\leq i\leq n$.
@@
$\mu_{i}$ $=$ $Y_{i}^{2}-\left[S_{i}+\eta\sum_{j=1,j\neq i}^{n+1}(Y_{i}-S_{i})\right],\ i=1,\ldots,n,$
$\mu_{n+1}$ $=$ $Y_{n+1}^{2}-\left[S_{n+1}+\eta\sum_{j=1}^{n}(Y_{i}-S_{i})\right].$
@@
$Y_{i}^{2}-\left[S_{i}+\eta\sum_{j=1}^{n}(Y_{i}-S_{i})-\eta(Y_{i}-S_{i})+\eta(Y_{n+1}-S_{n+1})\right]\geq Y_{n+1}^{2}-\left[S_{n+1}+\eta\sum_{j=1}^{n}(Y_{i}-S_{i})\right],$
which implies $\mu_{i}\geq\mu_{n+1}$ if and only if
$(Y_{n+1}+\eta/2)^{2}\leq\eta^{2}/4+\eta c_{i}+d_{i}.$ (7)
By the assumption $\eta\geq 2(\sqrt{\max\{0,c_{i}^{2}-d_{i}\}}-c_{i})$, we know the right-hand side of (7) is positive. Therefore, $\mu_{i}\geq\mu_{n+1}$ if and only if
$Y_{n+1}\in I_{i}=(-\sqrt{\eta^{2}/4+\eta c_{i}+d_{i}}-\eta/2,\sqrt{\eta^{2}/4+\eta c_{i}+d_{i}}-\eta/2)=(a_{i},b_{i}),\ i=1,\ldots,n.$
The same argument in the proof of Theorem 2 shows that $Y_{n+1}\in C_{\alpha}(X_{n+1},Z^{n})$ if and only if $\sum_{i=1}^{n}1_{\{\mu_{i}\geq\mu_{n+1}\}}>r=\lfloor(n+1)\alpha\rfloor$. Note that $Y_{n+1}\not\in(a_{(r_{1})},b_{(r_{3})})$ if and only if $Y_{n+1}$ does not belong to at most r1r-1 $I_{i}$’s, or equivalently, $Y_{n+1}\in(a_{(r_{1})},b_{(r_{3})})$ if and only if $Y_{n+1}$ belongs to at least $r$ $I_{i}$’s. Therefore, $Y_{n+1}\in C_{\alpha}(X_{n+1},Z^{n})$ if and only if $Y_{n+1}\in(a_{(r_{1})},b_{(r_{3})})$. ∎

### v2

@@ 2.1 Answers to questions regarding monotonicity of $M$
Answer to Question (V): No. Monotonicity of $M$ is not a necessary condition for the exact determination of $C_{\alpha}(X_{n+1},Z^{n})$.
@@
Remark. We did not ask the converse of Question (V), i.e., whether the monotonicity of $M$ implies that $C_{\alpha}(X_{n+1},Z^{n})$ can be determined exactly, because that question seems to be too broad to be well-defined. The next example illustrates this point.
@@
2.2 Questions regarding general properties of full conformal prediction
Answer to Question (VI): No. (4) is not a necessary condition for $C_{\alpha}(X_{n+1},Z^{n})$ to be an interval.
Example 7.
$C_{\alpha}(X_{n+1},Z^{n})=(-\infty,a_{(k)}),$
Answer to Question (VII): Yes. Monotonicity of the plausibility function $\mathsf{pl}_{x_{n+1},z^{n}}$ is a necessary condition for $C_{\alpha}(X_{n+1},Z^{n})$ to be a one-sided interval.
Proof.
Example 8.
Consider Example 7.
Example 9.
Consider Example 4.
@@ 3 Proposed strategy
If $\lfloor(n+1)\alpha\rfloor<1$, then the $(1-\alpha)\%$ prediction region $C_{\alpha}(X_{n+1},Z^{n})$ given by (2) is $\mathbb{R}$.
@@
Since $\lfloor(n+1)\alpha\rfloor<1$, we have $Y_{n+1}\in C_{\alpha}(X_{n+1},Z^{n})$ if and only if $\mathsf{pl}_{X_{n+1},Z^{n}}(Y_{n+1})>\lfloor(n+1)\alpha\rfloor/(n+1)$ if and only if $\sum_{i=1}^{n+1}1_{\{\mu_{i}\geq\mu_{n+1}\}}>\lfloor(n+1)\alpha\rfloor$. In view of the fact that $1_{\{\mu_{n+1}\geq\mu_{n+1}\}}=1$, we conclude that $C_{\alpha}(X_{n+1},Z^{n})=\mathbb{R}$. ∎
Thus, if we want a nontrivial conformal prediction region (i.e., $C_{\alpha}(X_{n+1},Z^{n})\neq\mathbb{R}$), we must require $\lfloor(n+1)\alpha\rfloor\geq 1$. For the remainder of this article, we let $r_{1}=\min\{n,\lfloor(n+1)(1-\alpha)\rfloor+1\}$, $r_{2}=(n+1)-\lfloor(n+1)(1-\alpha)\rfloor$, and $r_{3}=\max\{1,\lfloor(n+1)\alpha\rfloor\}$.
The negative answers to most questions in the previous section show that it is challenging to give a hard-and-fast rule for choosing a non-conformity measure so that we can overcome the aforementioned challenges. In the extant literature, many existing statistical models, such as ordinary linear regression, ridge regression, lasso, and kernel density estimation, have been used to create non-conformity measures ( e.g., Vovk et al. 2005; Lei et al. 2013, 2014; Lei 2019). As discussed in Section 1, in the full conformal prediction framework, hardly any non-conformity measure created using a complicated model allows us to both determine $C_{\alpha}(X_{n+1},Z^{n})$ exactly and control its shape. We uphold the philosophy that estimation and prediction warrant different approaches. For example, a wrong model can perform predictions well and even outperform the true data-generating model in some cases, though the same cannot be said for estimation (Shmueli 2010). We treat conformal prediction as a purely model-free method. Therefore, we do not use any statistical model to create a non-conformity measure. Our philosophy does not apply to split conformal prediction, where a complicated model is often employed to make the test absolute regression residuals small so that the split conformal prediction interval is tight. It is evident that if (4) holds, then we can address the three practical challenges simultaneously. There are numerous choices of the non-conformity measures that can lead to (4). Driven by the principle of parsimony, we prefer something simple, such as a linear polynomial in data, as in Hong (2026a). However, to obtain a two-sided/bounded prediction interval, we will need something other than (4). In fact, a multivariate quadratic polynomial of data will suffice, as we will see below. Intuitively, a higher order polynomial is likely to cause the solution to the inequality $\mu_{i}\geq\mu_{n+1}$ to be a union of disjoint intervals, rendering $C_{\alpha}(X_{n+1},Z^{n})$ to be a union of disjoint intervals too. This intuition is clear from the calculation in Example 6 and the proofs of Theorem 3 and Theorem 5. Therefore, we propose the following non-conformity measure.
@@
$M(B,z)=(\beta_{2}y^{2}+\beta_{1}y)+\left[\gamma\sum_{j=1}^{p}x_{j}+\sum_{i=1}^{n}\left(\eta y_{i}-\delta\sum_{j=1}^{p}x_{ij}\right)\right],$ (7)
where $\beta_{1},\beta_{2},\gamma$, $\delta$, and $\eta$ are constant parameters to be chosen at the discretion of the data scientist. Indeed, the main function of these parameters is to control the shape of the conformal prediction regions. For example, if $\beta_{2}=0$ and $\beta_{1}=1$, then the resulting conformal prediction region $C_{\alpha}(X_{n+1},Z^{n})$ will be a $(-\infty,a)$-shaped interval; if $\beta_{2}=0$ and $\beta_{1}=-1$, then $C_{\alpha}(X_{n+1},Z^{n})$ will be a $(a,\infty)$-shaped interval; when $\beta_{2}\neq 0$, $\beta_{1}=0$, $\eta=0$, and $\gamma=-1$, the resulting conformal prediction region will be a $(a,b)$-shaped interval. The parameters $\gamma$, $\eta$, and $\delta$ are the weights a data scientist attaches to the information in $Y_{1},\ldots,Y_{n}$, $X_{1},\ldots,X_{n}$, and $X_{n+1}$, respectively. The next three theorems show that this non-conformity measure enables us to address the aforementioned three challenges.
@@
Suppose $0<\alpha<1$, $\lfloor(n+1)\alpha\rfloor\geq 1$, and the non-conformity measure is given by (7). If $\beta_{2}=0$, $\beta_{1}=1$, $\gamma,\delta\in\mathbb{R}$, and $1-\eta>0$, then the $100(1-\alpha)\%$ conformal prediction region $C_{\alpha}(X_{n+1},Z^{n})$ is the one-sided interval $(-\infty,a_{(r_{1})})$, where $a_{i}=Y_{i}+\frac{(\delta+\gamma)}{1-\eta}\sum_{j=1}^{p}(X_{ij}-X_{(n+1)j})$ for $1\leq i\leq n$ and $a_{(k)}$ is the $k$-th ordered value of $a_{1},\ldots,a_{n}$.
@@
$\mu_{i}$ $=$ $M(Z^{n+1}\backslash\{Z_{i}\},Z_{i})=Y_{i}+\left[\gamma S_{i}+\sum_{j=1,j\neq i}^{n+1}(\eta Y_{j}-\delta S_{j})\right],\ i=1,\ldots,n,$
$\mu_{n+1}$ $=$ $M(Z^{n},Z_{i})=Y_{n+1}+\left[\gamma S_{n+1}+\sum_{j=1}^{n}(\eta Y_{j}-\delta S_{j})\right].$
@@
$(1-\eta)Y_{n+1}\leq(1-\eta)Y_{i}+(\gamma+\delta)(S_{i}-S_{n+1}).$
Since $1-\eta>0$, the last display implies $\mu_{i}\geq\mu_{n+1}$ if and only if $Y_{n+1}\leq Y_{i}+\frac{(\delta+\gamma)}{1-\eta}(S_{i}-S_{n+1})$ for $1\leq i\leq n$. Therefore, the theorem follows from (2) and the definition of the plausibility function $\mathsf{pl}_{x_{n+1},z^{n}}$. ∎
Remark. If we disregard the information of $X$ and rely on information of $Y$ only, we know $(-\infty,Y_{(r_{1})})$ is a finite-sample valid $100(1-\alpha)\%$ prediction interval for $Y$ (e.g., Wilks 1941; Fligner and Wolfe 1976). We believe the additional information provided by $X$ is generally useful, and is represented by the term $\frac{(\delta+\gamma)}{1-\eta}\sum_{j=1}^{p}(X_{ij}-X_{(n+1)j})$ in this theorem. Note that choosing the weight $\frac{(\delta+\gamma)}{1-\eta}<1$ will not necessarily shorten $C_{\alpha}(X_{n+1},Z^{n})$ in this case, since $\sum_{j=1}^{p}(X_{ij}-X_{(n+1)j})$ can be either positive or negative. A similar remark applies to the next two theorems.
@@
Suppose $0<\alpha<1$, $\lfloor(n+1)\alpha\rfloor\geq 1$, and the non-conformity measure is given by (7). If $\beta_{2}=0$, $\beta_{1}=-1$, $\gamma,\delta\in\mathbb{R}$, and $1-\eta<0$, then the $100(1-\alpha)\%$ conformal prediction region $C_{\alpha}(X_{n+1},Z^{n})$ is the one-sided interval $(a_{(r_{2})},\infty)$, where $a_{i}=Y_{i}+\frac{(\delta+\gamma)}{\eta-1}\sum_{j=1}^{p}(X_{ij}-X_{(n+1)j})$ for $1\leq i\leq n$.
@@
Suppose $0<\alpha<1$, $\lfloor(n+1)\alpha\rfloor\geq 1$, and the non-conformity measure is given by (7). If $\beta_{2}=1$, $\beta_{1}=0$, $\gamma=-1$, $\eta=0$, and $\delta$ satisfies the inequality
$1+\max_{i^{\prime}}\left\{Y_{i^{\prime}}^{2}/\sum_{j=1}^{p}(X_{(n+1)j}-X_{i^{\prime}j})\right\}\leq\delta\leq 1+\min_{i^{\prime\prime}}\left\{Y_{i^{\prime\prime}}^{2}/\sum_{j=1}^{p}(X_{(n+1)j}-X_{i^{\prime\prime}j})\right\},$ (8)
where $i^{\prime},i^{\prime\prime}\in{1,\ldots,n}$ such that $\sum_{j=1}^{p}(X_{(n+1)j}-X_{i^{\prime}j})>0$ and $\sum_{j=1}^{p}(X_{(n+1)j}-X_{i^{\prime\prime}j})<0$. Then the $100(1-\alpha)\%$ conformal prediction region $C_{\alpha}(X_{n+1},Z^{n})$ is the bounded equal-tailed interval $(-a_{(r_{3})},a_{(n-r_{3}+1)})$, where $a_{i}=\sqrt{Y_{i}^{2}+(1-\delta)\sum_{j=1}^{p}(X_{(n+1)j}-X_{ij})}$ for $1\leq i\leq n$.
@@
$\mu_{i}$ $=$ $Y_{i}^{2}-S_{i}-\delta\sum_{j=1,j\neq i}^{n+1}S_{j},\ i=1,\ldots,n,$
$\mu_{n+1}$ $=$ $Y_{n+1}^{2}-S_{n+1}-\delta\sum_{j=1}^{n}S_{j}.$
@@
$Y_{n+1}^{2}\leq Y_{i}^{2}+(1-\delta)(S_{n+1}-S_{i}).$ (9)
When $\delta$ satisfies (8), the right-hand side of (9) is positive. Therefore, $\mu_{i}\geq\mu_{n+1}$ if and only if $Y_{n+1}$ belongs to
$I_{i}=\left(-\sqrt{Y_{i}^{2}+(1-\delta)(S_{n+1}-S_{i})},\sqrt{Y_{i}^{2}+(1-\delta)(S_{n+1}-S_{i})}\right)=(-a_{i},a_{i}),\ i=1,\ldots,n.$
Now put $I_{(i)}=(-a_{(n-i+1)},a_{(i)})$ for $i=1,\ldots,n$. Then $I_{(1)}\subseteq I_{(2)}\subseteq\ldots\subseteq I_{(n)}$. The same argument in the proof of Theorem 2 shows that $Y_{n+1}\in C_{\alpha}(X_{n+1},Z^{n})$ if and only if $\sum_{i=1}^{n+1}1_{\{\mu_{i}\geq\mu_{n+1}\}}>\lfloor(n+1)\alpha\rfloor$, of equivalently, $\sum_{i=1}^{n}1_{\{\mu_{i}\geq\mu_{n+1}\}}\geq r_{3}$. Note that $Y_{n+1}\in(a_{(r_{3})},b_{(n-r_{3}+1)})$ if and only if $Y_{n+1}$ belongs to at least $r_{3}$ $I_{(i)}$’s. Therefore, $Y_{n+1}\in C_{\alpha}(X_{n+1},Z^{n})$ if and only if $Y_{n+1}\in(a_{(r_{3})},a_{(n-r_{3}+1)})$. ∎
Remark. The left-hand side of (8) is increasing in $n$ and the right-hand side of (8) is decreasing in $n$. If $\sum_{j=1}^{p}(X_{(n+1)j}-X_{ij})=0$ for all $i=1,\ldots,n$, we can let $\delta=1$. In practice, we can simply take $\delta=1+a\underset{i^{\prime\prime}}{\min}\left\{Y_{i^{\prime\prime}}^{2}/\sum_{j=1}^{p}(X_{(n+1)j}-X_{i^{\prime\prime}j})\right\}$, where $\sum_{j=1}^{p}(X_{(n+1)j}-X_{i^{\prime\prime}j})<0$ and $0<a<1$.
For a prediction interval, its mean length is a measure of efficiency. Our main goal is to circumvent the aforementioned challenges in the full conformal prediction framework, not to construct a nearly optimally efficient prediction interval, although the latter can be an important topic for future research. However, for any confidence level $1-\alpha$, there is no finite-sample valid $100(1-\alpha)\%$ prediction interval with the shortest mean length, as the next theorem shows. Therefore, one shall not aim at such an “ideal” prediction interval.
Theorem 6.
For a given sample size $n$ and any confidence level $0<1-\alpha<1$, there is no finite-sample valid (distribution-free) $100(1-\alpha)\%$ prediction interval with shortest mean length.
Proof.
Suppose the statement of the theorem is false. Then there will be a finite-valid $100(1-\alpha)\%$ prediction interval $(L(X_{n+1},Z^{n}),U(X_{n+1},Z^{n}))$ for some confidence level $0<1-\alpha<1$ such that $0<d_{n}=\mathsf{E}[U(X_{n+1},Z^{n}))-L(X_{n+1},Z^{n}))]<\infty$ is the shortest mean length of a finite-sample valid $100(1-\alpha)\%$ prediction interval, irrespective of the distribution of $Y$. Let $0<l<r<n+1$ such that $(r-l)/(n+1)\geq 1-\alpha$. Then $(Y_{(l)},Y_{(r)})$ is also a finite-sample valid $100(1-\alpha)\%$ prediction interval. We can always find $\sigma>0$ sufficiently small such that
$0<\frac{2(n-1)\sigma}{\sqrt{2n-1}}<d_{n}.$
Now consider the case where $Y$ follows a distribution such that $Var(Y)=\sigma^{2}$. By the bounds of $E(Y_{(1)})$ and $E(Y_{(n)})$ (e.g., David and Nagara 2003, Section 4.2), we have
$0<\mathsf{E}\left[Y_{(r)}-Y_{(l)}\right]\leq\mathsf{E}\left[Y_{(n)}-Y_{(1)}\right]\leq\frac{2(n-1)\sigma}{\sqrt{2n-1}}<d_{n},$
contradicting the definition of $d_{n}$. ∎